Starting with an expression for U(S.V), show that m(V) = (dU/dV)T is given by Tt(v)= (dp/dT)V-P.
Answer #1

s= entropy, U= intern al energy, = volume, T= Temperature and P= Pressure. U=U CSV) – 0 du = arus les + ular - © From first lon dq=Tds Putting the value of dq and dw in the above equation ③ we get du= Tds -Pdv - 4 dividing by dv and taking T as a con

Then from equation ③ we get Av=MT= Coule = 7 (2P) - P/